what is the time period of oscillation for rigid hinged body
1
Expert's answer
2020-09-03T14:10:31-0400

solution:-

consider a rigid body of mass (M) hinged at point R as shown in figure.there is no friction between hinge and body.



if body is displaced from equilibrium by an angle "\theta" . then restoring torque can be written as


"\tau=-Mgl\sin\theta" .....eq.1

where "l" is the distance from hinged point R to center of gravity.

for small oscillation

"\sin\theta≈\theta"

and torque for a rigid body can be given


"\tau=I\alpha" ............eq.2

where "I" is moment of inertia of body about hinged point and "\alpha" is angular acceleration of body.

by the equation 1 and 2 we can written as


"-Mgl\theta=I\alpha"

"\frac{\theta}{\alpha}=-\frac{I}{Mgl}"

there are "-\theta" is proportional to "\alpha" so its simple harmonic motion. and time period can be given as


"T=2\pi\sqrt\vert\frac{\theta}{\alpha}\vert"


"\fcolorbox{green}{yellow}{$T=2\pi\sqrt\frac{I}{Mgl}$}"

this is the time period of oscillation for rigid hinged body.